OpenStudy (anonymous):

During rush hour, buses leave Main Street every seven minutes. The first part of the schedule is shown below. Main Street -7:41-7:48-7:55-8:02 Rush hour ends at 9:00 A.M. Leon gets to the bus stop at 8:25. At what time will Leon's bus depart at Main Street? Show Your Work: Answer: Explain, how you found your answer

OpenStudy (anonymous):

8:25 - 8:02 = 23 min 23 min / 7 = 3 ( 2 min remainder ) 7 - 2 = 5 mins => 8: 25 + 5 min = 8: 30 AM is depart time.